OpenStudy (jchick):

Carlos solved the system of equations below. What mistake did he make in his work? 2x + y = 5 x − 2y = 10 y = 5 − 2x x − 2(5 − 2x) = 10 x + 10 + 4x = 10 5x + 10 = 10 5x = 0 x = 0 2(0) + y = 5 y = 5

OpenStudy (jchick):

a. He should have substituted 5 + 2x b. He did not distribute correctly c. He combined like terms incorrectly, it should have been 4x instead of 5x d. He made no mistake

OpenStudy (jchick):

y really equals -3

OpenStudy (anonymous):

hey texaschic101 can u finish helping me please

OpenStudy (texaschic101):

first we will solve for y in the 1st equation 2x + y = 5 y = -2x + 5 or y = 5 - 2x now we sub 5 - 2x in for y in the 2nd equation x - 2y = 10 x - 2(5 - 2x) = 10 -- now we distribute through the parenthesis x - 10 + 4x = 10 (MISTAKE HERE....distributed wrong

OpenStudy (anonymous):

2x + y = 5 --> y = -2x + 5 x - 2y = 10 x - 2(-2x + 5)= 10 x + 4x - 10= 10 5x - 10 = 10 5x = 20 x = 4 Just plug four into your original equation for x and you'll get the y value